# Decremental $(1+\epsilon)$-Approximate Maximum Eigenvector: Dynamic Power Method

### v1

@@ Abstract
We present a dynamic algorithm for maintaining $(1+\epsilon)$-approximate maximum eigenvector and eigenvalue of a positive semi-definite matrix $A$ undergoing decreasing updates, i.e., updates which may only decrease eigenvalues. Given a vector $v$ updating $A\leftarrow A-vv^{\top}$, our algorithm takes $\tilde{O}(\mathrm{nnz}(v))$ amortized update time, i.e., polylogarithmic per non-zeros in the update vector. Our technique is based on a novel analysis of the influential power method in the dynamic setting. The two previous sets of techniques have the following drawbacks (1) algebraic techniques can maintain exact solutions but their update time is at least polynomial per non-zeros, and (2) sketching techniques admit polylogarithmic update time but suffer from a crude additive approximation. Our algorithm exploits an oblivious adversary. Interestingly, we show that any algorithm with polylogarithmic update time per non-zeros that works against an adaptive adversary and satisfies an additional natural property would imply a breakthrough for checking psd-ness of matrices in $\tilde{O}(n^{2})$ time, instead of $O(n^{\omega})$ time.
@@ 1 Introduction
The study of the change in the spectrum of a matrix undergoing updates spans over five decades, with [Gol73] providing the first algebraic characterization of the change for positive semi-definite matrices via the secular equation. This result offered an explicit formula for exactly computing all new eigenvalues when a matrix undergoes a single rank-one update, assuming knowledge of the entire eigen-decomposition of the initial matrix. Subsequently, [BNS78] showed how to additionally compute eigenvectors explicitly, and handle the case of repeated eigenvalues of the initial matrix. A decade later, [AGG88] extended the works of [Gol73, BNS78] to compute all new eigenvalues and eigenvectors of a positive semi-definite matrix undergoing an update of the form of a small-rank matrix, or a single batch update. Further works extend these techniques to computing singular values [Sta08], as well as computing new eigenvalues and eigenvectors when only the partial eigen-decomposition of the initial matrix is known [MSS19]. However, all these works require a full eigen-decomposition of the initial matrix and only handle a single update to the matrix.
@@
Now consider the simpler task of only maintaining the maximum eigenvalue and eigenvector of a matrix $\bm{\mathit{A}}$ as it undergoes updates. As we have seen above, known methods based on algebraic techniques require full spectral information before computing the new eigenvalues, and maintaining the entire eigen-decomposition can be slow. Sparsification-based algorithms are fast but only achieve large additive approximations of at least $\epsilon n$, which is not quite satisfactory. Works on streaming PCA, which have a similar goal of maintaining large eigenvectors focus on optimizing the space complexity instead of runtimes [AL17]. Previous works based on dynamically maintaining the matrix inverse can maintain $(1+\epsilon)$-multiplicative approximations to the maximum eigenvalues of an $n\times n$ symmetric matrix undergoing single-entry updates with an update time of $O(n^{1.447}/\mathrm{poly}(\epsilon))$ [San04]. This was further improved to $O(n^{1.407}/\mathrm{poly}(\epsilon))$[BNS19]. The update time is even slower for row, column, or rank-one updates. In any case, it takes polynomial time per non-zeros of the updates. This leads to the following natural question:
@@
In this paper, we study the problem of maintaining a $(1+\epsilon)$-multiplicative approximation to the maximum eigenvalue and eigenvector of a positive semi-definite matrix as it undergoes a sequence of rank-one updates that may only decrease the eigenvalues. We now formally state our problem.
@@
[…]
@@
[…]
@@ 1.1 Our Results
Our algorithm is a novel adaptation of the classical power method [TB22] to the dynamic setting, along with a new analysis that may be of independent interest.
Our work can also be viewed as the first step towards generalizing the dynamic algorithms for solving positive linear programs of [BKS23] to solving dynamic positive semi-definite programs. We discuss this connection in detail in Section 5.
@@ 1.2 Towards Separation between Oblivious and Adaptive adversaries
We first state the barrier formally, and then state our result for adaptive adversaries. Recall that, given a matrix $\bm{\mathit{A}}$, the condition number of $\bm{\mathit{A}}$ is $\frac{\max_{x\mathrel{\mathop{:}}\textrm{unit}}\|\bm{\mathit{A}}\bm{\mathit{x}}\|}{\min_{x\mathrel{\mathop{:}}\textrm{unit}}\|\bm{\mathit{A}}\bm{\mathit{x}}\|}$.
@@
Recall that $A$ is a psd matrix iff there exists $\bm{\mathit{X}}$ such that $\bm{\mathit{A}}=\bm{\mathit{X}}\bm{\mathit{X}}^{\top}$. The matrix of $\bm{\mathit{X}}$ is called a vector realization of $\bm{\mathit{A}}$, and note that $\bm{\mathit{X}}$ is not unique. The problem above asks to compute a (highly accurate) vector realization of $\bm{\mathit{A}}$. Both eigendecomposition and Cholesky decomposition of $\bm{\mathit{A}}$ give a solution for 1.3.222Given an eigendecomposition $\bm{\mathit{A}}=\bm{\mathit{Q}}\Lambda\bm{\mathit{Q}}^{\top}$ where $\Lambda$ is a non-negative diagonal matrix and $\bm{\mathit{Q}}$ is an orthogonal matrix, we set $\bm{\mathit{X}}=\bm{\mathit{Q}}\Lambda^{1/2}$. Given a Cholesky decomposition $\bm{\mathit{A}}=\bm{\mathit{L}}\bm{\mathit{L}}^{\top}$ where $\bm{\mathit{L}}$ is a lower triangular matrix, we set $\bm{\mathit{X}}=\bm{\mathit{L}}$. Banks et al [Ban+22] showed how to compute with high accuracy an eigendecomposition of a psd matrix in $O(n^{\omega+\eta})$ time for any constant $\eta>0$, where $\omega>2.37$ is the matrix multiplication exponent. Hence, a vector realization of $\bm{\mathit{A}}$ can be found in the same time. Observe that, when $\delta<\kappa$, 1.3 is at least as hard as certifying that a given matrix $\bm{\mathit{A}}$ is a psd matrix. It is a notorious open problem whether certifying the psd-ness of a matrix can be done faster than $o(n^{\omega})$ even when the condition number is polynomial in $n$. Therefore, we view 1.3 as a significant barrier and formalize it as follows.
@@
Our result states that if there is an algorithm for Problem 1.1 against adaptive adversaries such that the algorithm returns $\bm{\mathit{w}}_{t}$ satisfying an additional property, which we state below, then Hypothesis 1.4 is false.
@@
Proposition 3.17 is quite natural. Consider a “bad” eigenvector $\bm{\mathit{u}}_{i}$ whose corresponding eigenvalue is very small, i.e., less than half of the maximum one. It states that the projection of $\bm{\mathit{w}}_{t}$ along $\bm{\mathit{u}}_{i}$ must be very small, i.e., a polynomial factor smaller than the projection along the maximum eigenvector. Indeed, the power method returns an approximate maximum eigenvector that strongly satisfies 1.5 (see Lemma 3.3). Our algorithm, however, returns $\bm{\mathit{w}}_{t}$’s that only satisfy this property at certain snapshots and not at every step.
We now state our hardness result, which shows that any algorithm for 1.1 that works against adaptive adversaries cannot have sub-polynomial update time without breaking the barrier above.
@@
Assuming 1.4, there is no algorithm for 1.1 that finds $\bm{\mathit{w}}_{t}$’s additionally satisfying 1.5, and given parameters $n$ and $\epsilon=\min\mathopen{}\mathclose{{}\left\{1-\frac{1}{n^{o(1)}},\frac{1-\delta}{1+\delta}}\right\}$ as input, works against an adaptive adversary in time $n^{o(1)}\cdot\mathopen{}\mathclose{{}\left(nnz(\bm{\mathit{A}}_{0})+\sum_{t=1}^{T}nnz(\bm{\mathit{v}}_{i})}\right)$.
Understanding the power and limitations of dynamic algorithms against an oblivious adversary vs. an adaptive adversary has become one of the main research programs in the area of dynamic algorithms. An algorithm is said to be robust if it works against an adaptive adversary. Recently, Beimel et al. [Bei+22] took a complexity-theoretic approach to understanding the gap between the two models of adversaries. From the upper bound side, they showed the first black-box transformation that can transform any non-robust dynamic algorithm for any estimation problem [^3] In estimation problems, an algorithm needs to output one number for each query, such as the size of minimum cut or the distance of any given pair of vertices. to the corresponding robust algorithm with some overhead. From the lower bound side, they also gave the first separation by showing the two problems that admit fast non-robust algorithms, but any robust algorithms require polynomially slower update time under a strong cryptographic assumption. Further work by Bateni et al. [Bat+23] shows a separation between the adversaries for the $k$-center clustering problem, however, their results are based on the existence of a black box which the adaptive adversary can control. Therefore, finding a natural dynamic problem that separates oblivious and adaptive adversaries is still a wide-open problem. We believe this paper makes progress towards this goal, and if one could show the lower bound without requiring 1.5, we would have a solution to this long-studied open problem.
@@
In the following sections, we begin with some preliminaries required to prove our results in Section 2, followed by our algorithm and its analysis in Section 3, and our conditional lower bound in Section 4. In Section 5 we show connections with positive semi-definite programs, and finally, in Section 6, we present some open problems.
@@ 2 Preliminaries
The vector $\bm{\mathit{v}}$ has entries that are from $N(0,1)$. Now, every $\bm{\mathit{v}}_{i}^{2}$ follows a $\chi^{2}$ distribution. From Lemma 1 of [LM00] we have the following tail bound for a sum of $\chi^{2}$ random variables,
@@ 3 Algorithms against an Oblivious Adversary
[…]
@@
Algorithm 3 Update algorithm at time $t$ ($A_{t-1},r_{t},\bm{\mathit{W}}_{t-1}=[w^{(r)}_{t-1}\mathrel{\mathop{:}}r=1,\cdots R],\epsilon$ are maintained)
@@ 3.1 Proof Overview
Let $\epsilon>0$ and let $\bm{\mathit{W}}=[\bm{\mathit{w}}^{(1)},\dots,\bm{\mathit{w}}^{(R)}]$ be as defined in Line 9 in the execution of PowerMethod($\epsilon,\bm{\mathit{A}}$). Let $\bm{\mathit{v}}_{1},\cdots,\bm{\mathit{v}}_{k}$ be a sequence of updates and define $\bm{\widetilde{\mathit{A}}}=\bm{\mathit{A}}-\sum_{i=1}^{k}\bm{\mathit{v}}_{i}\bm{\mathit{v}}_{i}^{\top}$. Suppose that $\lambda_{\max}(\bm{\mathit{A}})\geq 1-\epsilon$ and $\bm{\mathit{w}}^{\top}\bm{\widetilde{\mathit{A}}}\bm{\mathit{w}}<1-40\epsilon$ for all $\bm{\mathit{w}}\in\bm{\mathit{W}}$. Then, with probability at least $1-\frac{50\log\frac{n}{\epsilon}}{n^{2}}$, for some $\nu\in\mathcal{I}$,
@@
We prove this lemma in Section 3.4. Intuitively speaking, it means that, whenever Line 4 of Algorithm 3 is executed, there is some important level $\nu$ such that an $\Omega(\epsilon/\mathrm{poly}\log(n/\epsilon))$-fraction of eigenvalues of $\bm{\mathit{A}}$ at level $\nu$ have decreased in value.
@@
The following lemma shows that the projection of $\bm{\mathit{w}}$ on $T_{\nu}-{\tilde{\mathit{T}}}$ is small when $\mathrm{dim}(T_{\nu}-{\tilde{\mathit{T}}})\mathrel{\mathop{:}}=q_{\nu}$ is roughly at most an $\epsilon$-factor of $\mathrm{dim}(T_{\nu})\mathrel{\mathop{:}}=d_{\nu}$, and $q_{\nu}$ is still at least logarithmic. We will use this lemma to characterize the projection of $\bm{\mathit{w}}$ on $T_{\nu}$ for $\nu\in\mathcal{I}$. Recall that $\Pi_{\nu}$ is a projection matrix that projects any vector to the space $T_{\nu}-{\tilde{\mathit{T}}}$.
@@ 4 Conditional Lower Bounds for an Adaptive Adversary
In this section, we will prove a conditional hardness result for algorithms against adaptive adversaries. In particular, we will prove Theorem 1.6. Before proving our result, let us consider Algorithm 4 for solving 1.3. The only step in Algorithm 4 whose implementation is not specified is Line 8. We will implement this step using an algorithm for 1.1.
@@
11:     if $0\leq\mu_{T}\leq\frac{(1+\epsilon)\mu_{1}\delta}{\kappa}$ then
12:         return $\bm{\mathit{X}}=\frac{1}{\sqrt{10}}\begin{bmatrix}\sqrt{\mu_{1}}\bm{\mathit{w}}_{1}&\sqrt{\mu_{2}}\bm{\mathit{w}}_{2}&\cdots&\sqrt{\mu_{T}}\bm{\mathit{w}}_{T}\end{bmatrix}$
13:     else
14:         return False:$\bm{\mathit{A}}$ is not PSD
We first prove the guarantees of the output of Algorithm 4 when $\bm{\mathit{w}}_{t}$’s satisfy 1.5 for all $t$.
@@
Next, if $\bm{\mathit{A}}$ is not psd, then we will prove that $\mu_{T}\geq\frac{(1-\epsilon)\mu_{1}}{\kappa}.$ This implies that $\mu_{T}>\frac{(1+\epsilon)\mu_{1}\delta}{\kappa}$ from the value of $\epsilon$. and so the algorithm will return False by the condition at Line 11. Since $\bm{\mathit{A}}$ is not psd, there exists a unit vector $\bm{\mathit{v}}$ such that $\bm{\mathit{v}}^{\top}\bm{\mathit{A}}\bm{\mathit{v}}<0$. We can lower bound $\mu_{T}$ as follows.
$\mu_{T}\geq(1-\epsilon)\|\bm{\mathit{A}}_{T}\|\geq(1-\epsilon)|\bm{\mathit{v}}^{\top}\bm{\mathit{A}}_{T}\bm{\mathit{v}}|.$
@@
where the first inequality is because $\bm{\mathit{v}}^{\top}\bm{\mathit{A}}_{T}\bm{\mathit{v}}=\bm{\mathit{v}}^{\top}\bm{\mathit{A}}\bm{\mathit{v}}-\|\bm{\mathit{X}}^{\top}\bm{\mathit{v}}\|^{2}<\bm{\mathit{v}}^{\top}\bm{\mathit{A}}\bm{\mathit{v}}<0$. Combining the two inequalities, we get $\mu_{T}\geq\frac{(1-\epsilon)\mu_{1}}{\kappa}$, which concludes the proof. ∎

### v2

@@ Abstract
We present a dynamic algorithm for maintaining $(1+\epsilon)$-approximate maximum eigenvector and eigenvalue of a positive semi-definite matrix $A$ undergoing decreasing updates, i.e., updates which may only decrease eigenvalues. Given a vector $v$ updating $A\leftarrow A-vv^{\top}$, our algorithm takes $\tilde{O}(\mathrm{nnz}(v))$ amortized update time, i.e., polylogarithmic per non-zeros in the update vector.
Our technique is based on a novel analysis of the influential power method in the dynamic setting. The two previous sets of techniques have the following drawbacks (1) algebraic techniques can maintain exact solutions but their update time is at least polynomial per non-zeros, and (2) sketching techniques admit polylogarithmic update time but suffer from a crude additive approximation.
Our algorithm exploits an oblivious adversary. Interestingly, we show that any algorithm with polylogarithmic update time per non-zeros that works against an adaptive adversary and satisfies an additional natural property would imply a breakthrough for checking psd-ness of matrices in $\tilde{O}(n^{2})$ time, instead of $O(n^{\omega})$ time.
@@ 1 Introduction
[…]
Now consider the simpler task of only maintaining the maximum eigenvalue and eigenvector of a matrix $\bm{\mathit{A}}$ as it undergoes updates. As we have seen above, known methods based on algebraic techniques require full spectral information before computing the new eigenvalues, and maintaining the entire eigen-decomposition can be slow. Sparsification-based algorithms are fast but only achieve large additive approximations of at least $\epsilon n$, which is not quite satisfactory. Works on streaming PCA, which have a similar goal of maintaining large eigenvectors focus on optimizing the space complexity instead of runtimes [allen2017first]. Previous works based on dynamically maintaining the matrix inverse can maintain $(1+\epsilon)$-multiplicative approximations to the maximum eigenvalues of an $n\times n$ symmetric matrix undergoing single-entry updates with an update time of $O(n^{1.447}/\mathrm{poly}(\epsilon))$ [sankowski2004dynamic]. This was further improved to $O(n^{1.407}/\mathrm{poly}(\epsilon))$[van2019dynamic]. The update time is even slower for row, column, or rank-one updates. In any case, it takes polynomial time per non-zeros of the updates. This leads to the following natural question:
@@
In this paper, we study the problem of maintaining a $(1+\epsilon)$-multiplicative approximation to the maximum eigenvalue and eigenvector of a positive semi-definite matrix as it undergoes a sequence of rank-one updates that may only decrease the eigenvalues. We note that this is equivalent to finding the minimum eigenvalue and eigenvector of a positive semi-definite matrix undergoing rank-one updates that may only increase the eigenvalues. We now formally state our problem.
@@
$\max_{\bm{\mathit{u}}:\textrm{unit}}\bm{\mathit{u}}^{\top}\bm{\mathit{A}}_{t}\bm{\mathit{u}}\geq\lambda_{t}\geq(1-\epsilon)\max_{\bm{\mathit{u}}:\textrm{unit}}\bm{\mathit{u}}^{\top}\bm{\mathit{A}}_{t}\bm{\mathit{u}},$ (1)
@@
$\bm{\mathit{w}}_{t}^{\top}\bm{\mathit{A}}_{t}\bm{\mathit{w}}_{t}\geq(1-\epsilon)\max_{\bm{\mathit{u}}:\textrm{unit}}\bm{\mathit{u}}^{\top}\bm{\mathit{A}}_{t}\bm{\mathit{u}}.$ (2)
@@ 1.1 Our Results
Our algorithm is a novel adaptation of the classical power method (see, e.g., [trefethen2022numerical]) to the dynamic setting, along with a new analysis that may be of independent interest.
Our work can also be viewed as the first step towards generalizing the dynamic algorithms for solving positive linear programs of [bhattacharya2023dynamic] to solving dynamic positive semi-definite programs. We discuss this connection in detail in Appendix A.
@@ 1.2 Towards Separation between Oblivious and Adaptive adversaries
We first state the barrier formally, and then state our result for adaptive adversaries. Recall that, given a matrix $\bm{\mathit{A}}$, the condition number of $\bm{\mathit{A}}$ is $\frac{\max_{x:\textrm{unit}}\|\bm{\mathit{A}}\bm{\mathit{x}}\|}{\min_{x:\textrm{unit}}\|\bm{\mathit{A}}\bm{\mathit{x}}\|}$.
@@
Recall that $A$ is a psd matrix iff there exists $\bm{\mathit{X}}$ such that $\bm{\mathit{A}}=\bm{\mathit{X}}\bm{\mathit{X}}^{\top}$. The matrix of $\bm{\mathit{X}}$ is called a vector realization of $\bm{\mathit{A}}$, and note that $\bm{\mathit{X}}$ is not unique. The problem above asks to compute a (highly accurate) vector realization of $\bm{\mathit{A}}$. Both eigendecomposition and Cholesky decomposition of $\bm{\mathit{A}}$ give a solution for 1.3.222Given an eigendecomposition $\bm{\mathit{A}}=\bm{\mathit{Q}}\Lambda\bm{\mathit{Q}}^{\top}$ where $\Lambda$ is a non-negative diagonal matrix and $\bm{\mathit{Q}}$ is an orthogonal matrix, we set $\bm{\mathit{X}}=\bm{\mathit{Q}}\Lambda^{1/2}$. Given a Cholesky decomposition $\bm{\mathit{A}}=\bm{\mathit{L}}\bm{\mathit{L}}^{\top}$ where $\bm{\mathit{L}}$ is a lower triangular matrix, we set $\bm{\mathit{X}}=\bm{\mathit{L}}$. Banks et al [banks2022pseudospectral] showed how to compute with high accuracy an eigendecomposition of a psd matrix in $O(n^{\omega+\eta})$ time for any constant $\eta>0$, where $\omega>2.37$ is the matrix multiplication exponent. Hence, a vector realization of $\bm{\mathit{A}}$ can be found in the same time. Observe that, when $\delta<\kappa$, 1.3 is at least as hard as certifying that a given matrix $\bm{\mathit{A}}$ is a psd matrix. It is a notorious open problem whether certifying the psd-ness of a matrix can be done faster than $o(n^{\omega})$ even when the condition number is polynomial in $n$. Therefore, we view 1.3 as a significant barrier and formalize it as follows.
@@
Our negative result states that, assuming Hypothesis 1.4, there is no algorithm for Problem 1.1 against adaptive adversaries with sub-polynomial update time that maintains $\bm{\mathit{w}}_{t}$ satisfying an additional property stated below.
@@
Assuming 1.4, there is no algorithm for 1.1 that maintains $\bm{\mathit{w}}_{t}$’s additionally satisfying 1.5, and given parameters $n$ and $\epsilon=\min\mathopen{}\mathclose{{}\left\{1-\frac{1}{n^{o(1)}},\frac{1-\delta}{1+\delta}}\right\}$ as input, works against an adaptive adversary in time $n^{o(1)}\cdot\mathopen{}\mathclose{{}\left(nnz(\bm{\mathit{A}}_{0})+\sum_{t=1}^{T}nnz(\bm{\mathit{v}}_{i})}\right)$.
Let us motivate 1.5 from several aspects below. First, in the static setting, this property can be easily satisfied since the static power method strongly satisfies it (see Lemma 3.3). Second, the statement of the property itself is a natural property that we might expect from an algorithm. Consider a “bad” eigenvector $\bm{\mathit{u}}_{i}$ whose corresponding eigenvalue is very small, i.e., less than half of the maximum one. It states that the projection of the output $\bm{\mathit{w}}_{t}$ along $\bm{\mathit{u}}_{i}$ should be very small, i.e., a polynomial factor smaller than the projection along the maximum eigenvector. This is intuitively useful because we do not want the output $\bm{\mathit{w}}_{t}$ to direct to the “bad” direction. It should mainly direct along the approximately maximum eigenvector. Third, we formalize this intuition and crucially exploit 1.5 to prove a reduction in Theorem 1.6. Specifically, 1.5 allows us to decrease eigenvalues of a psd matrix while maintaining the psd-ness, which is crucial for us. See Section 4 for more details. Lastly, our current algorithm from Theorem 1.2, a dynamic version of the power method, actually maintains $\bm{\mathit{w}}_{t}$’s that satisfy this property for certain snapshots, but not at every step. Unfortunately, we do not see how to strengthen our algorithm to satisfy 1.5 nor how to remove it from the requirement of Theorem 1.6. We leave both possibilities as open problems.
Understanding the power and limitations of dynamic algorithms against an oblivious adversary vs. an adaptive adversary has become one of the main research programs in the area of dynamic algorithms. However, finding a natural dynamic problem that separates oblivious and adaptive adversaries is still a wide-open problem.333 Beimel et al. [beimel2022dynamic] gave the first separations for artificial problems assuming a strong cryptographic assumption. Bateni et al. [bateni2023optimal] shows a separation between the adversaries for the $k$-center clustering problem, however, their results are based on the existence of a black box which the adaptive adversary can control. In this paper, we suggest the problem of maintaining approximate maximum eigenvectors as a natural candidate for the separation and give some preliminary evidence for this.
@@
In the following sections, we begin with some preliminaries required to prove our results in Section 2, followed by our algorithm and its analysis in Section 3, and our conditional lower bound in Section 4. In Section A we show connections with positive semi-definite programs, and finally, in Section 5, we present some open problems.
@@ 2 Preliminaries
The vector $\bm{\mathit{v}}$ has entries that are from $N(0,1)$. Now, every $\bm{\mathit{v}}_{i}^{2}$ follows a $\chi^{2}$ distribution. From Lemma 1 of [laurent2000adaptive] we have the following tail bound for a sum of $\chi^{2}$ random variables,
@@ 3 Algorithms against an Oblivious Adversary
$\|\bm{\mathit{w}}_{t}\|=1\quad\text{and}\quad\bm{\mathit{w}}_{t}^{\top}\bm{\mathit{A}}_{t}\bm{\mathit{w}}_{t}\geq 1-40\epsilon,$
@@
Algorithm 3 Update algorithm at time $t$ ($A_{t-1},r_{t},\bm{\mathit{W}}_{t-1}=[w^{(r)}_{t-1}:r=1,\cdots R],\epsilon$ are maintained)
@@ 3.1 Proof Overview
Let $\epsilon>0$ and let $\bm{\mathit{W}}=[\bm{\mathit{w}}^{(1)},\dots,\bm{\mathit{w}}^{(R)}]$ be as defined in Line 9 in the execution of PowerMethod($\epsilon,\bm{\mathit{A}}$). Let $\bm{\mathit{v}}_{1},\cdots,\bm{\mathit{v}}_{k}$ be a sequence of updates generated by an oblivious adversary and define $\bm{\widetilde{\mathit{A}}}=\bm{\mathit{A}}-\sum_{i=1}^{k}\bm{\mathit{v}}_{i}\bm{\mathit{v}}_{i}^{\top}$.
Suppose that $\lambda_{\max}(\bm{\mathit{A}})\geq 1-\epsilon$ and $\bm{\mathit{w}}^{\top}\bm{\widetilde{\mathit{A}}}\bm{\mathit{w}}<1-40\epsilon$ for all $\bm{\mathit{w}}\in\bm{\mathit{W}}$. Then, with probability at least $1-\frac{50\log\frac{n}{\epsilon}}{n^{2}}$, for some $\nu\in\mathcal{I}$,
@@
We prove this lemma in Section 3.4. Intuitively speaking, it means that, whenever Line 4 of Algorithm 3 is executed, there is some important level $\nu$ such that an $\Omega(\epsilon/\mathrm{poly}\log(n/\epsilon))$-fraction of eigenvalues of $\bm{\mathit{A}}$ at level $\nu$ have decreased in value. This is the crucial place where we exploit an oblivious adversary.
@@
The following lemma shows that the projection of $\bm{\mathit{w}}$ on $T_{\nu}-{\tilde{\mathit{T}}}$ is small when $\mathrm{dim}(T_{\nu}-{\tilde{\mathit{T}}}):=q_{\nu}$ is roughly at most an $\epsilon$-factor of $\mathrm{dim}(T_{\nu}):=d_{\nu}$, and $q_{\nu}$ is still at least logarithmic. We will use this lemma to characterize the projection of $\bm{\mathit{w}}$ on $T_{\nu}$ for $\nu\in\mathcal{I}$. Recall that $\Pi_{\nu}$ is a projection matrix that projects any vector to the space $T_{\nu}-{\tilde{\mathit{T}}}$.
@@ 4 Conditional Lower Bounds for an Adaptive Adversary
In this section, we will prove a conditional hardness result for algorithms against adaptive adversaries. In particular, we will prove Theorem 1.6. Consider Algorithm 4 for solving 1.3. The only step in Algorithm 4 whose implementation is not specified is Line 8. We will implement this step using an algorithm for 1.1.
@@
11:     $\sigma^{2}\leftarrow$PowerMethod($\epsilon,\bm{\mathit{A}}_{T}^{\top}\bm{\mathit{A}}_{T}$)
12:     if $0\leq\sigma\leq\frac{(1+\epsilon)\mu_{1}\delta}{\kappa}$ then
13:         return $\bm{\mathit{X}}=\frac{1}{\sqrt{10}}\begin{bmatrix}\sqrt{\mu_{1}}\bm{\mathit{w}}_{1}&\sqrt{\mu_{2}}\bm{\mathit{w}}_{2}&\cdots&\sqrt{\mu_{T}}\bm{\mathit{w}}_{T}\end{bmatrix}$
14:     else
15:         return False:$\bm{\mathit{A}}$ is not PSD
High-level idea.
Overall for our hardness result, we use the idea that an adaptive adversary can use the maximum eigenvectors returned to perform an update. This can happen $n$ times and in the process, we would recover the entire eigen-decomposition of the matrix, which is hard. Now consider Algorithm 4. We claim that Algorithm 4 solves 1.3. At the first glance, this claim looks suspicious because the input matrix for 1.3 might not be PSD, but the dynamic algorithm for 1.1 at Line 8 has any guarantees only when the matrices remain PSD. However, the reduction does work by crucially exploiting 1.5. The high-level idea is as follows.
  • If the input matrix $\bm{\mathit{A}}$ is initially PSD, then we can show that $\bm{\mathit{A}}_{t}$ remains PSD for all $t$ by exploiting 1.5, (see Lemma 4.1). So, the approximation guarantee of the algorithm at Line 8 is valid at all steps. From this guarantee, $\|\bm{\mathit{A}}_{T}\|$ must be tiny since we keep decreasing the approximately maximum eigenvalues (see Lemma 4.2). At the end, the reduction will return $\bm{\mathit{X}}$.
  • If the input matrix $\bm{\mathit{A}}$ is initially not PSD, there must exist a direction $\bm{\mathit{v}}$ such that $\bm{\mathit{v}}^{\top}\bm{\mathit{A}}\bm{\mathit{v}}<0$. Since in the reduction, we update $\bm{\mathit{A}}_{T}=\bm{\mathit{A}}-\bm{\mathit{W}}$ for some $\bm{\mathit{W}}\succeq 0$, we must have that $\bm{\mathit{v}}^{\top}\bm{\mathit{A}}_{T}\bm{\mathit{v}}<\bm{\mathit{v}}^{\top}\bm{\mathit{A}}\bm{\mathit{v}}$. That is, this negative direction remains negative or gets even more negative. It does not matter at all what guarantees the algorithm at Line 8 has. We still have that $\|\bm{\mathit{A}}_{T}\|$ cannot be tiny. We can distinguish whether $\|\bm{\mathit{A}}_{T}\|$ is tiny or not using the static power method at Line 11, and, hence, we will return False in this case (see Lemma 4.2).
Analysis.
We prove the guarantees of the output of Algorithm 4 when $\bm{\mathit{w}}_{t}$’s satisfy 1.5 for all $t$.
@@
We would like to point out that our parameter $\epsilon$ is quite large. This just implies that our reduction can work even if we find crude approximations to the maximum eigenvector as long as this is along the directions with large eigenvalue, since $\bm{\mathit{w}}$ also has to satisfy 1.5.
@@
Next, if $\bm{\mathit{A}}$ is not psd, then we will prove that $\sigma\geq\frac{(1-\epsilon)\mu_{1}}{\kappa}.$ This implies that $\sigma>\frac{(1+\epsilon)\mu_{1}\delta}{\kappa}$ from the value of $\epsilon$. and so the algorithm will return False by the condition at Line 12. Since $\bm{\mathit{A}}$ is not psd, there exists a unit vector $\bm{\mathit{v}}$ such that $\bm{\mathit{v}}^{\top}\bm{\mathit{A}}\bm{\mathit{v}}<0$. We can lower bound $\sigma$ as follows.
$\sigma\geq(1-\epsilon)\|\bm{\mathit{A}}_{T}\|\geq(1-\epsilon)|\bm{\mathit{v}}^{\top}\bm{\mathit{A}}_{T}\bm{\mathit{v}}|.$
@@
where the first inequality is because $\bm{\mathit{v}}^{\top}\bm{\mathit{A}}_{T}\bm{\mathit{v}}=\bm{\mathit{v}}^{\top}\bm{\mathit{A}}\bm{\mathit{v}}-\|\bm{\mathit{X}}^{\top}\bm{\mathit{v}}\|^{2}<\bm{\mathit{v}}^{\top}\bm{\mathit{A}}\bm{\mathit{v}}<0$. Combining the two inequalities, we get $\sigma\geq\frac{(1-\epsilon)\mu_{1}}{\kappa}$, which concludes the proof. ∎